OpenStudy (anonymous):

Solve using substitution: -y − 5x = -7 2y − 2x = 14

OpenStudy (anonymous):

get y by itself, and substitute into the other equation

OpenStudy (anonymous):

do you understand what im talking about?

OpenStudy (anonymous):

mm a little. so add y?

OpenStudy (anonymous):

add y and add 7 to both sides of the first equation

OpenStudy (anonymous):

so it has no solution?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

its multiple choice

OpenStudy (anonymous):

you end up with y=-5x+7 correct

OpenStudy (anonymous):

Infinitely many solutions No solution (7/5, 14) (0, 7)

OpenStudy (anonymous):

lets work out the problem so take that and substitute it into the second equation to get 2(-5x+7) − 2x = 14

OpenStudy (anonymous):

distribute to get -10x+14-2x=14 -12x=0

OpenStudy (anonymous):

so your x value is 0 now plug 0 into x for the first equation to find y

OpenStudy (anonymous):

so y=-5(0)+7 y coordinate is 7

OpenStudy (anonymous):

giving you (0,7)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

thankss.!!! :)

OpenStudy (anonymous):

np you understand?

OpenStudy (anonymous):

yeah i get this. i have an answer for the next one but idk if its right..can you help me with my next one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve using substitution: -2x − y + 1 = 0 2x = -13 − 3y

OpenStudy (anonymous):

ok so lets get y by itself again

OpenStudy (anonymous):

add

OpenStudy (anonymous):

i got 4,-7

OpenStudy (anonymous):

2(x)=-13-3(-2x+1) 2x=-13+6x-3 4x=16 x=4

OpenStudy (anonymous):

y=-2(4)+1 y=-8+1 y=-7

OpenStudy (anonymous):

(4,-7)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np glad to help

OpenStudy (anonymous):

:)